# EVA: An Embodied World Model for Future Video Anticipation

### v1

@@ Abstract
World models integrate raw data from various modalities—such as images and language to simulate comprehensive interactions in the world, thereby displaying crucial roles in fields like mixed reality and robotics. Yet, applying the world model for accurate video prediction is quite challenging due to the complex and dynamic intentions of the various scenes in practice. In this paper, inspired by the human rethinking process, we decompose the complex video prediction into four meta-tasks that enable the world model to handle this issue in a more fine-grained manner. Alongside these tasks, we introduce a new benchmark named Embodied Video Anticipation Benchmark (EVA-Bench) to provide a well-rounded evaluation. EVA-Bench focused on evaluating the video prediction ability of human and robot actions, presenting significant challenges for both the language model and the generation model. Targeting embodied video prediction, we propose the Embodied Video Anticipator (EVA), a unified framework aiming at video understanding and generation. EVA integrates a video generation model with a visual language model, effectively combining reasoning capabilities with high-quality generation. Moreover, to enhance the generalization of our framework, we tailor-designed a multi-stage pretraining paradigm that adaptatively ensembles LoRA to produce high-fidelity results. Extensive experiments on EVA-Bench highlight the potential of EVA to significantly improve performance in embodied scenes, paving the way for large-scale pre-trained models in real-world prediction tasks.
[…]
@@ 1 Introduction
A world model integrates raw data from various modalities such as images and language, to imagine how the world evolves as an agent behaves (Ha & Schmidhuber, 2018). It aims to understand the physical world they encounter and generate high-fidelity videos, allowing embodied agents to plan, evaluate, and simulate operations in their own neural space. Such a model shows great potential, particularly in fields like mixed reality, autonomous driving (Gao et al., 2024; Wang et al., 2023b), gaming (Bruce et al., 2024), and robotics (Yang et al., 2023).
Imaging a world model with future prediction ability. It would not only understand environments but also interact by anticipating future actions and generating predictive videos. Such predictive videos could serve as interactive guidelines, akin to a product manual in mixed reality, a driving instructor, or a robot’s planning imagination (Du et al., 2023a), greatly enhancing the decision-making process and helping the embodied agents towards Artificial General Intelligence.
This paper aims to enable embodied agents to generate predictive videos using multimodal instructions, allowing them to visualize potential outcomes before taking action.
However, existing methods (Zhou et al., 2024) usually focus on conditional simulation and overlook the complexities of multimodal and multi-level time-scale predictions. Moreover, benchmarking such prediction tasks involves several critical steps, data collection, model training, validation, and continuous refinement, which remain largely unexplored. To briefly, there are two main challenges in this task:
Figure 2: Giving the instruction and observation as input, the world model planning can be decomposed into several meta-questions. With the combination of these meta-questions, Rethinking, the world model gives a proper output.
  • 1)
    Lack of Benchmark: Video generation involves producing frame-level actions based on previous motion cues, guided by high-level language instructions (Xing et al., 2024). However, benchmarking these multilevel and multimodal predictions poses challenges due to the complexity of achieving a unified understanding of coarse-grained high-level language instructions.
  • 2)
    Training Models: Recent pre-trained general models usually struggle with significant inaccuracies and illusion outputs in embodied scenes (Xing et al., 2023). In contrast, the task-specific models exhibit limited generalization capabilities (Yan et al., 2021; Zhou et al., 2024). A major challenge lies in generating high-fidelity action predictions while preserving a comprehensive understanding of the world within a unified framework, particularly when dealing with complex instructions.
To benchmark this task, we start by reformulating the problem in a coarse-to-fine manner. Given common sense that frame-level prediction of video generation is a minimum step of prediction, it is not enough for a task-level prediction. As shown in Figure 2, each task-level prediction step requires a comprehensive understanding of the current situation, followed by a next-step prediction of what will happen. Such prediction, in the human mind, could further become pictures or so-called imagination. Then we wonder whether this imagination satisfied this task or not. Such a rethinking process keeps going until this prediction step is completed. Therefore, with the inspiration of recent work (Wei et al., 2022; Liu et al., 2024b; Kawaharazuka et al., 2024), we address these challenges by defining the task within embodied scenes into four meta tasks, Action-Description, How-To, Finish-Thinking, Next-Step, as shown in Figure 1. Furthermore, to evaluate such cases, we introduce an embodied video anticipation benchmark (EVA-Bench) to evaluate the video prediction task.
We further propose a novel world model, the Embodied Video Anticipator (EVA), which is capable of generating future prediction videos for embodied scenes, thereby enhancing the interaction experience between humans and machines. To achieve this, we re-collect published datasets from multiple domains to create a diverse dataset and introduce a multi-stage pretraining method. We build an embodied video predictor through a video generation model(VDM) and visual language model(VLM). Additionally, we implement a cross-attention alignment model for end-to-end optimization. We introduce a few-shot adaptation method with an Ensemble of LoRA Hu et al. (2021) to ensure the model can adapt to specific cases efficiently and incrementally.
@@
  • Human-Rethinking-Inspired Rethinking Generation: The paper formulates the problem of video prediction by defining it within embodied scenes into four meta tasks: Action-Description, How-To, Finish-Thinking, and Next-Step. This structured approach helps in addressing the complexities of multi-level and multimodal predictions.
  • Embodied Video Anticipator: We formulate the Embodied Video Anticipator (EVA) framework, a multimodal understanding and generation model for embodied scene generation. We integrate multi-stage pretraining, cross-attention alignment, and few-shot adaptation with an Ensamble-LoRA. Our model can give consistent and long-horizon video predictions.
  • Embodied Video Prediction Benchmark Creation: We introduce EVA-Bench, a comprehensive benchmark designed to evaluate the performance of world models in predicting future events within embodied scenes. This benchmark provides a standardized way to measure and compare different models, addressing the challenge of defining and evaluating video anticipation.
@@ 2 Related Work
Video Generation With the advent of diffusion-based visual generation models, there has been significant progress in extending the capabilities of video generation. For instance, models like VideoCrafter (Chen et al., 2023; 2024) and VideoPoet (Kondratyuk et al., 2023) have demonstrated impressive abilities in generating high-quality video. Moreover, video generation models with image conditions like Dynamicrafter (Xing et al., 2023), Stable Video Diffusion (Blattmann et al., 2023) and Animatediff (Guo et al., 2023) meet impressive generation quality and have already been used in many areas. Such weakness also happens in some long video generation methods (Wang et al., 2023a; Yin et al., 2023). These video-generation models lack reasoning abilities and still struggle with consistency and understanding, so recent works also combine generation models with LLMs (He et al., 2024b).
@@
World Model World models aim to provide future predictions based on current observations. This concept has been explored in various domains, including Genie (Bruce et al., 2024), which shows interesting ability in gaming simulation, Vista (Gao et al., 2024; Wang et al., 2023b), etc., in autonomous driving. Video prediction is a special world-model-like task, Seer (Gu et al., 2024), AID (Xing et al., 2024) adapting image-to-video generation model to predict the motion of future frames. Additionally, world models such as RoboDreamer (Zhou et al., 2024) and AVDC (Ko et al., 2023) have been utilized as robot simulators. Unisim (Yang et al., 2023), for instance, combining pre-trained web-scale data with embodied videos expands world models’ applications. VLP (Du et al., 2023b) integrated language and video generation models for robot planning but stayed at the concept level. A task-level video predictor is still needed.
Embodied Dataset and Benchmark On the other hand, such multimodal benchmarks are still missing. Open-X (Padalkar et al., 2023) provides comprehensive robot data, Ego-exo4d (Grauman et al., 2024) provides multi-level language annotation to key steps, and Webvid-10M-Action (Xing et al., 2023) updates the action annotation to the internet video dataset. The MMWorld (He et al., 2024a) benchmark includes embodied tasks for visual-question-answer tasks. However, there is still a lack of a video-language-to-video-language dataset that can define and benchmark the prediction ability of the world model.
In summary, world models, whether as simulators or future predictors, require a well-defined task benchmark and model paradigm to advance the field.
@@
Figure 3: self-ask (Press et al., 2022) inference pipeline of EVA. Given the visual observation and human questions as input, EVA would first generate fixed frames of videos and related text answers. Then, the model prompts itself to check the task completion status; if the predicted video is not finished, EVA keeps generating the extended frames until the task completion judgment is true.
3 Task Formulation and Benchmark
To benchmark the video prediction task, we first format it. In this section, we will first give a comprehensive task formulation explanation in Section 3.1, and then we will describe how we turn it into four meta-generation problems and how we benchmark them in Section 3.2.
3.1 Task Formulation
To begin with, we noticed that there is an existing multi-level of language prediction (Cheng et al., 2024); in this paper, we define frame-level, task-level, high-level, etc. For a simple example, task-level prediction is direct language instruction, like ”pick up the book”, and the serial task-level formulate a higher level instruction, e.g., ”Cleaning the table.” Modern video generation models can make frame-level predictions by extending the previous frames.
This paper aims to solve the video prediction problem at the task level. Given common sense that frame-level prediction of video generation is a minimum step of prediction, it is not enough for a task-level prediction. So, the formulation of a task-level prediction and a multi-round of frame-level prediction are required.
In our model, we define the world model function $\mathcal{M}$ as taking visual observations $O$ and a question $Q$ as inputs, producing a predicted video $\hat{V}$ and a text response $\hat{A}$ as outputs:
$(\hat{V},\hat{A})=\mathcal{M}(O,Q)$ (1)
To ensure reliable task-level predictions, the model first comprehends the observed video $O$. It then decides whether to provide frame-level or task-level predictions. This decision is transformed into predicted videos $\hat{V}$, which are then iteratively refined until they meet task-level requirements. This process involves four key tasks: Action-Description (demonstrating understanding), How-To (indicating simulation quality), Finish-Thinking (checking task-level generation satisfaction), and Next-Step (showing task-level prediction capability). The world model can make further tasks or higher-level predictions only with a solid foundation in these four tasks.
3.2 Embodied Video Anticipator Benchmark
We establish a benchmark for task-level video prediction. Initially, we present a comprehensive design of evaluation metrics that evaluate the video-language to video-language generation task. We have named this the EVA-Score (EVAS). The representation of EVAS is as follows:
$\textit{EVAS}_{\text{normalized}}^{t}=\frac{\sum_{i=1}^{n_{t}}\alpha_{t,i}\psi_{t,i}+\sum_{j=1}^{m_{t}}\beta_{t,i}\iota_{t,j}}{n_{t}+m_{t}}$ (2)
@@ 3.2 Embodied Video Anticipator Benchmark
Therefore, we describe their language and video metrics for each task as follows:
Action-Description We expect the text prompt to be short and simple, containing five key elements: subject, verb, object, location, and destination, following the basic English paradigm. We first evaluate this task using a series of text metrics to determine whether the output description closely matches the ground truth. We also added GPT4 (OpenAI, 2024) metric to judge the keywords and CLIP (Radford et al., 2021) score to evaluate the description quality.
Finsh-Thinking This task aims to judge whether the frame-level prediction has completed the task. Real-world task generation has dynamic time durations based on different tasks. However, current video generation models can only generate a limited number of frames. In the VQA task, the finish-thinking output “Yes” or “No” could be used to calculate the accuracy. During generation, apart from the existing metric, we put forward a Goal Completion Estimation (GCE) metric to compare the last frame of ground truth and the generation frame.
How-To For a “How-To” question, giving the instruction, the world model should be able to turn it into visual output. We introduce several video metrics to formulate the video evaluation metrics, including Motion Smoothness and Multi-stage Consistency inspired by VBench Huang et al. (2024). Combined with the text metrics, these form the correlation metric that represents the quality of “How-To” tasks. The final EVA-Score could benchmark the quality of How-To.
Next-Step The last and most important task is to evaluate the task-level prediction of the next step in the video prediction system. We reorganized high-level labeling into task-level Next-Step QA problems to formulate the dataset. While evaluating, we compared the predicted action. We predicted video with the ground truth dataset and obtained the cross-modality correlation among these features, which is also included in our EVA-Score.
These questions help decompose the task into more specific and actionable components, enabling the model to better understand and predict future events across various scenarios. We anticipate that this benchmark and dataset will serve as a foundational setting for evaluating the capabilities of world models. We include the detailed dataset and benchmark information in the Appendix A.5.
@@ 4 Embodied Video Anticipator
To train the EVA, we designed a multimodal generative large language model, including two main pre-trained models for multimodal prediction. As we described in Section 3, we formulate the problem as a video-language to video-language task, which can be represented as equation 1. EVA uses a multistage training strategy, designed an Ensemble method for domain-specific LoRA Hu et al. (2021) for VDM, interaction tokens to achieve this complex visual prediction task. We describe these key elements in detail in the following section.
4.1 VLM Backbone
VLM Backbone In EVA, we utilize a 7B Visual Language Model (VLM) backbone called ChatUniVi (Jin et al., 2024). The visual foundation model is CLIP (ViT-L/14) (Radford et al., 2021), while the LLM backbone is the Vicuna-v1.5 (Zheng et al., 2023) model. Moreover, an adaptive parameter-free token clustering algorithm  (Jin et al., 2024) is used to reduce the number of video tokens. This backbone setting introduces additional computational overhead and allows training to remain within the 2048-token limit. We first trained the visual project on the Video-Descrtption dataset. We then independently trained the full parameter fine-tuning of LLM on the QA Instruction Tuning dataset, as shown in Figure 4.
4.2 VDM Backbone
We utilize a 1.5B pre-trained Latent video diffusion model (VDM) called Dynamicrafter (Xing et al., 2023) for the generation model. The VDM conditions on image, fps, and language embedding features to generate a 16-frame video in 2 seconds. It UNet incorporates temporal and spatial transformers, as illustrated in Figure 4, and is pre-trained on a large-scale web video dataset, enabling it to generate dynamic content for open-domain images.
4.3 Ensemble of LoRA
$\vspace{-0.5em}W_{d}=W+\Delta W_{d}=W+A_{d}B_{d}^{T}$ (3)
$\vspace{-0.5em}g_{d}=\text{softmax}(f(\text{Task Token}))$ (4)
$\vspace{-0.5em}\hat{W}=W+\sum_{d}g_{d}\Delta W_{d}$ (5)
4.4 Cross-Attention Generation Adapter

### v2

# Empowering World Models with Reflection for Embodied Video Prediction

@@ Abstract
Video generation models have made significant progress in simulating future states, showcasing their potential as world simulators in embodied scenarios. However, existing models often lack robust understanding, limiting their ability to perform multi-step predictions or handle Out-of-Distribution (OOD) scenarios. To address this challenge, we propose the Reflection of Generation (RoG), a set of intermediate reasoning strategies designed to enhance video prediction. It leverages the complementary strengths of pre-trained vision-language and video generation models, enabling them to function as a world model in embodied scenarios. To support RoG, we introduce Embodied Video Anticipation Benchmark(EVA-Bench), a comprehensive benchmark that evaluates embodied world models across diverse tasks and scenarios, utilizing both in-domain and OOD datasets. Building on this foundation, we devise a world model, Embodied Video Anticipator (EVA), that follows a multistage training paradigm to generate high-fidelity video frames and apply an autoregressive strategy to enable adaptive generalization for longer video sequences. Extensive experiments demonstrate the efficacy of EVA in various downstream tasks like video generation and robotics, thereby paving the way for large-scale pre-trained models in real-world video prediction applications. The video demos are available at https://sites.google.com/view/icml-eva.
Machine Learning, ICML
Figure 1: The illustration of the Reflection of Generation(RoG). Giving the instruction and observation as input, the world model gives a proper output with the combination of understanding, prediction, and generation.
@@ 1 Introduction
The rapid development of video generation technologies has brought predictive models to the forefront of building embodied world simulators (Agarwal et al., 2025), where the ability to anticipate future status is critical in embodied scenarios (Wang et al., 2024b). These generated predictive videos can serve as interactive guidelines, mixed-reality product manuals, driving instructors (Gao et al., 2024; Wang et al., 2023b), gaming assistants (Bruce et al., 2024), or even a robot’s planning imagination (Du et al., 2023a; Yang et al., 2023). By imagining how the world evolves as an agent behaves, this prediction capability significantly enhances decision-making in embodied scenarios, providing a foundation for robust world models (Ha & Schmidhuber, 2018).
However, existing video generation models for robotics (Zhou et al., 2024; Yang et al., 2023) primarily focus on conditional simulation, neglecting the critical role of understanding both current and predicted states. Therefore, these models are limited to generating fixed-frame sequences and single-round predictions. Moreover, they lack the ability to detect flawed outputs, relying heavily on human intervention. These limitations severely degrade their performance in Out-of-Distribution (OOD) scenarios and undermine the fundamental goal of an embodied world simulator.
To address these challenges, we propose Reflection-of-Generation (RoG), inspired by recent advances in reasoning and alignment (Wei et al., 2022; Liu et al., 2024b; Kawaharazuka et al., 2024). As shown in  Figure 1, RoG integrates intermediate reasoning steps into the video generation process, enabling iterative self-correction and deeper understanding. It separates the understanding and generation tasks, allowing for a modular and interpretable design. Therefore, RoG can combine any Visual Language Model (VLM) and Video Generation Model (VDM) into a unified world model. By introducing reflection as a critical step, RoG empowers the world model to iteratively refine its predictions, and generate longer video sequences.
To facilitate the training and evaluation of RoG, we introduce EVA-Bench, a comprehensive benchmark designed for embodied video anticipation. EVA-Bench decompose the RoG within embodied scenes into four meta-tasks: Action-Description, How-To, Finish-Thinking, Next-Step, as illustrated in  Figure 2. These meta-tasks simplify data collection and evaluation by focusing on both understanding and generation aspects. EVA-Bench includes both in-domain and out-of-distribution videos, leveraging diverse datasets from egocentric and robotics scenarios. It evaluates models using a combination of Video Question-Answering (VQA) metrics and pixel-level video prediction metrics.
Building on this benchmark, we propose EVA (Embodied Video Anticipator), a unified understanding and generation model tailored to the RoG. EVA employs a chunk-wise autoregressive paradigm, integrating reflection conditions to adaptively extend the length of generated video sequences. By supervising the generation process with VLMs, EVA achieves consistent and self-correcting video generation, enhancing the interaction between humans and machines in embodied scenarios.
@@
  • Reflection-of-Generation(RoG): We propose RoG, a novel strategy that integrates intermediate reasoning steps into video generation, enabling self-correction and deeper understanding in embodied scenarios.
  • EVA-Benchmark: We create EVA-Bench, a benchmark that evaluates world models across diverse embodied tasks with in-domain and OOD scenarios, using both understanding and generation metrics.
  • Embodied Video Aiticipator: We design EVA, a unified understanding and generation world model that generates longer, more consistent videos through chunk-wise autoregressive reflection.
Extensive experiments on EVA-Bench highlight the strong performance of RoG in both in-domain and OOD tasks. Furthermore, to validate the applicability of EVA in robot planning, we evaluate the model using a robot simulator (Mees et al., 2022; Brohan et al., 2022), demonstrating that EVA and RoG effectively support real-world task execution.
@@ 2 Related Work
Video Generation With the advent of diffusion-based visual generation models, there has been significant progress in extending the capabilities of video generation. For instance, models like VideoCrafter (Chen et al., 2023, 2024) and VideoPoet (Kondratyuk et al., 2023) have demonstrated impressive abilities in generating high-quality video. Moreover, video generation models with image conditions like Dynamicrafter (Xing et al., 2023), Stable Video Diffusion (Blattmann et al., 2023) and Animatediff (Guo et al., 2023) meet impressive generation quality and have already been used in many areas. Such weakness also happens in some long video generation methods (Wang et al., 2023a; Yin et al., 2023). These video-generation models lack reasoning abilities and still struggle with consistency and understanding, so recent works also combine generation models with LLMs (He et al., 2024).
@@
Embodied World Model World models aim to provide future predictions based on current observations. This concept has been explored in various domains, including Genie (Bruce et al., 2024), which shows interesting ability in gaming simulation, Vista (Gao et al., 2024; Wang et al., 2023b), etc., in autonomous driving. Video prediction is a special world-model-like task, Seer (Gu et al., 2024), AID (Xing et al., 2024) adapting image-to-video generation model to predict the motion of future frames. Additionally, world models such as RoboDreamer (Zhou et al., 2024), Enerverse (Huang et al., 2025) and AVDC (Ko et al., 2023) have been utilized video generation as robot simulators. Unisim (Yang et al., 2023), for instance, combining pre-trained web-scale data with embodied videos expands world models’ applications. VLP (Du et al., 2023b) integrated language and video generation models for robot planning but stayed at the concept level. A task-level video predictor is still needed.
@@
Figure 3: RoG inference pipeline of EVA with chunk-wise frame extension. Given the visual observation and human questions as input, EVA would first generate fixed frames of videos and related text answers. Then, the model prompts itself to check the task completion status; if the predicted video is not finished, EVA keeps generating the extended frames until the task completion judgment is true.
3 Reflection-of-Generation
In this section, we introduce the RoG by first defining hierarchical levels of visual prediction and then the basic process of the RoG world model.
3.1 Hierarchy of Video Prediction
The video prediction task, begin with an initial observation $O_{0}$ at time step 0 and a task instruction $I$, is categorized into hierarchical levels of complexity: frame-level prediction, task-level prediction, and serial task-level prediction.
Frame-Level Prediction: The objective at this level is to generate subsequent video frames $O_{t}(t\geq 1)$, conditioned on the initial observation $O_{0}$. This formulates a next-frame prediction problem, where the model approximates the distribution: $P(O_{t}|O_{0},I),\quad t=1,2,\dots$
Task-Level Prediction: Task-level prediction focuses on achieving a high-level instruction $I$, such as “pick up the book”. The goal is to model the sequence of predictions $\{O_{t}\}$ required to complete the task: $P(\{O_{t}\}_{t=1}^{T}|O_{0},I)$ Here, $T$ represents the time horizon necessary to complete the task. This level abstracts away low-level frame transitions, instead emphasizing coherent task execution.
Serial Task-Level Prediction: The more complex level involves sequentially completing multiple sub-tasks to fulfill a broader goal. For instance, instruct “clean the table”, the model must decompose it into sub-tasks (e.g., “remove objects”, “wipe the surface”) and execute them sequentially.
Task-Level Prediction via a Unified World Model: This paper focuses on task-level prediction as the core of the video prediction task, bridging frame-level generation and task reasoning. The central objective of task-level prediction is to simulate the state of the world, which is the role of the world model, denoted by $\mathcal{W}$. Formally:
$\mathcal{W}(O_{0},I)\approx P(\{O_{t}\}_{t=1}^{T}|O_{0},I)$ (1)
The world model processes $O_{0}$ and $I$, dynamically combining frame generation and task reasoning to generate the sequence $\{O_{t}\}$. The predictions are refined iteratively until the task requirements are met. This hierarchical approach ensures that task-level prediction incorporates both low-level frame synthesis and high-level task satisfaction, addressing the demands of complex video prediction tasks.
3.2 Reflection of Generation World Model
The RoG world model introduces a dynamic feedback mechanism that integrates an understanding module and a generation module to iteratively refine task-level predictions, as shown in  Figure 3. The model is defined as follows:
$\mathcal{W}_{\text{RoG}}(O_{0},I)=\text{Reflect}(H)\cdot P(\{O_{t}\}_{t=1}^{T}|O_{0},I,H)$ (2)
Where $\{O_{t}\}_{t=1}^{T}$ is a sequence of predicted observations, $H$ is the encoded hidden state representing the prediction status, $\text{Reflect}()$ is the understanding module, which provides three types of responses: extend, regenerate, or output. The whole inference process is as shown in  Algorithm 1. The algorithm iteratively refines a prediction sequence $\{O_{t}\}$ using a reflection mechanism. It begins by encoding the initial observation $O_{0}$, task instruction $I$, and current predictions into a high-dimensional representation $H$.
Based on $H$, the reflection mechanism determines whether to extend the sequence by generating new frames, regenerate flawed predictions, or finalize the sequence if it is complete. This process continues until the predictions converge or meet the task’s stopping criteria, thereby ensuring adaptive and accurate video generation in embodied scenarios.
Algorithm 1 Reflection of Generation World Model ($\mathcal{W}_{\text{RoG}}$)
  Input: Initial observation $O_{0}$, task instruction $I$
  Output: Sequence of predictions $\{O_{t}\}_{t=1}^{T}$
  Initialize: Prediction sequence $\{O_{t}\}\leftarrow\emptyset$
  repeat
     $H\leftarrow\text{Encode}(O_{0},I,\{O_{t}\})$ Understanding Module encodes input states
     $\hat{V}\leftarrow\text{Reflect}(H)$ Generate reflection output based on $H$
     if $\hat{V}=\text{Extend}(\{O_{t}\})$ then
        Extend the prediction sequence $\{O_{t}\}$
     else if $\hat{V}=\text{Regenerate}(\{O_{t}\})$ then
        Regenerate the prediction sequence $\{O_{t}\}$
     else if $\hat{V}=\text{Output}(\{O_{t}\})$ then
        Finalize and output the prediction sequence $\{O_{t}\}$
        break
     end if
  until Prediction sequence converges or breaks
3.3 RoG World Model Task Decomposition
The RoG world model is evaluated through four sub-tasks with corresponding metrics, decomposing video prediction into actionable components for a comprehensive assessment. The tasks are as follows:
@@ 3.3 RoG World Model Task Decomposition
Finish-Thinking: This task assesses whether the frame-level prediction should be extended to complete the task. Metrics include the accuracy of binary outputs (”Yes”/”No”). We introduce the Goal Completion Estimation (GCE) to compare the generated frame with the ground truth, Fréchet Video Distance (Unterthiner et al., 2018) multiple video generation quality metrics from VBench (Huang et al., 2024).
How-To: This task transforms task instructions into visual outputs and text responses, evaluated by video metrics and text metrics. The combined score of language and video forms the EVA-Score for How-To tasks.
Next-Step: Predict the next action in video sequences and language instructions. The metrics are the same as for How-To, including the accuracy of the QA at the task level and the correlation between the intermodality between the predicted frames and the ground truth.
@@ 4 Embodied Video Anticipator
As we formulate the problem as a video prediction task as equation 2, we propose the world model EVA. EVA includes two main pre-trained models for multimodal prediction, uses a multistage training strategy, and uses an Ensemble method for domain-specific LoRA (Hu et al., 2021) for VDM and interaction tokens to achieve this complex visual prediction task. We describe these key elements in detail in the following section.
4.1 Model Structure
EVA integrates a 7B VLM backbone and a 1.5B VDM to enable high-quality video generation for diverse tasks. The VLM backbone combines a CLIP (ViT-L/14) visual encoder with the Vicuna-v1.5 language model, using a parameter-free token clustering algorithm to reduce video token overhead. Special interaction tokens, such as $<image>$ and $<IMG_{P}>$, bridge visual and textual inputs, aligning with task-specific processing. The VDM generates 16-frame videos in 2 seconds by conditioning on image, FPS, and language embeddings, leveraging temporal and spatial transformers pre-trained on large-scale video datasets. To enhance adaptability, the Ensemble-LoRA method trains low-rank adaptation modules for each domain (e.g., human-egocentric, real-robot). It combines them using a task token-controlled gating system, ensuring efficient multi-domain generalization. Additionally, a cross-attention generation adapter aligns VLM hidden features with VDM text embeddings via linear transformation, optimized using diffusion denoising loss for superior video generation quality. Detailed module structure is described in  Appendix A.